The crew of an enemy spacecraft attempts to escape from your spacecraft by moving away from you at 0.283 of the speed of light. But all is not lost! You launch a space torpedo toward the foe at 0.341 of the speed of light with respect to you. At what speed in kilometers per second does the enemy crew observe the torpedo approaching its spacecraft?
Answer #1

Velocity, u' = (u - v)/(1 - uv/c^2)

= (0.341 - 0.283)c/(1 - (0.341 x 0.283))

= (0.058c)/(0.9034)

= 0.058 x 3 x 10^8/0.9034

= 1.93 x 10^7 m/s = 1.93 x 10^4 km/s

Comment in case any doubt please rate my answer ...